OpenStudy (anonymous):

if you have (1,-2) and (-3,1) how do you find the slope and y-intercept

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

i found the slope it is -3/4 and i put it into y-y1=m(x-x1) and i cant get the y -intercept

hartnn (hartnn):

to get 'y' intercept from any equation, you just put x =0 in that equation :)

OpenStudy (anonymous):

ok so the question on my worksheet says" explain how you would use the points (1,-2) and (-3,1) to graph a linear equation and find the slope and y-inercept

hartnn (hartnn):

so you could graph it right ? and y-intercept from graph is even easier, its where you line crosses x-axis

hartnn (hartnn):

i mean crosses, y-axis

OpenStudy (anonymous):

okay thanks.

hartnn (hartnn):

welcome ^_^ and since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

haha thanks:)